# Tomoya Shiota,∗,†,a,b Kenji Ishihara,†,b and Wataru Mizukami∗,a,b

### v2

@@
[Uncaptioned image] Computational screening is indispensable for the efficient design of high-entropy alloys (HEAs), which hold considerable potential for catalytic applications. However, the chemical space of HEAs is exponentially vast with respect to the number of constituent elements, making even machine learning-based screening calculations time-intensive. To address this challenge, we propose a rapid method for predicting HEA properties using data from monometallic systems (or few-component alloys). Central to our approach is the newly introduced local surface energy (LSE) descriptor, which captures local surface reactivity at atomic resolution. By applying linear regression, we successfully screened adsorption energies of molecules on HEAs based on LSEs derived from monometallic systems. Furthermore, we developed high-precision models by employing both classical and quantum machine learning. Our method enabled CO adsorption-energy calculations for 1000 quinary nanoparticles, comprising 201 atoms each, within a few days, considerably faster than density functional theory, which would require hundreds of years or neural network potentials, which would have taken hundreds of days. The proposed approach accelerates the exploration of the vast HEA chemical space, facilitating the design of novel catalysts.
@@ 1 Introduction
Fig.  1: (a) Near-exponential increase in the number of nanoparticles (NPs) and adsorption sites as the number of elements in the NPs increases from monometallic to multicomponent systems. (b) Relationship between the molecular adsorption energy and the newly introduced descriptor, i.e., the local surface energy (LSE), for monometallic and high-entropy alloy (HEA) NP systems. The LSE descriptor developed in this study effectively captures the variation in adsorption energy across different adsorption sites and NP compositions. (c) Fast catalytic property prediction workflow using neural newtwork potentials (NNPs). The adsorption energy ${E_{\text{ad}}}$ is predicted from the LSE of the HEA NP surface prior to adsorption. (d) Efficient construction of the prediction model for CO adsorption energy using density functional theory (DFT) or NNPs, with the LSE descriptor. The LSE is computed by NNPs. The model is built through linear regression.
@@
To circumvent these limitations, neural network potentials (NNPs) based on the Behler–Parrinello framework 25, 26, 27, 28, 29 and graph neural networks 30, 31, 32, 33, offer promising solutions. Universal NNPs can encompass extensive elemental diversity and achieve high computational efficiency while maintaining accuracy on par with that of DFT 31, 32, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46. Recently, NNPs specializing in HEAs have emerged, made more lightweight through knowledge distillation 47. These advances have accelerated the prediction of catalytic properties; however, computational challenges remain.
In contrast, descriptor-based machine learning models offer scalability by predicting adsorption energies through generalized coordination numbers, d-band centers, and local atomic environments, bypassing direct energy assessments 48, 49, 50, 11, 51, 52. In addition to the coordination-number descriptor, models have been proposed for predicting the adsorption energies of the remaining candidates by regressing the adsorption energies obtained from several first-principles calculations 10. Nonetheless, their applicability to HEAs is hampered by the complexity of alloy compositions and their dependence on extensive first-principles calculations.
@@ 2.1 Target systems and problems
We made the following assumptions regarding the structure and composition of HEA NPs: the structure is a truncated octahedron NP with 201 atoms in a face-centered cubic (fcc) arrangement (see 1(c)), the elemental composition ratio is as uniform as possible, and the atomic arrangement is randomly determined following a uniform distribution. We also assumed that the molecule occupied only a single top site. In HEA201, there are 122 sites. Even with these assumptions, the number of adsorption sites on HEA NPs is approximately $1.3\times 10^{138}$ (Figure 1(a)).
As a specific demonstration system, we measured the adsorption energies of CO molecules on IrPdPtRhRu NPs. In 2020, Wu et al. successfully synthesized IrPdPtRhRu NPs with nearly identical experimental composition ratios  6. The investigation of the adsorption characteristics of CO molecules is useful for evaluating the catalytic properties of HEAs 11, 53. As precursor systems for developing predictive models for HEAs, the monometallic NPs are of the truncated octahedron type corresponding to $\mathrm{M}_{\mathrm{n}}$ n=38, 79, 116, 201. For the on-top adsorption of the CO molecule on monometallic NPs $\mathrm{M}_{201}$, only the irreducible adsorption sites are calculated, as represented by the black line in Figure 1(d). The adsorption energy $E_{\mathrm{ad}}$ is calculated as follows:
@@ 2.2 Local surface energy (LSE) descriptor
where $E_{\mathrm{atom}}^{\mathrm{surf}}$ and $E_{\mathrm{atom}}^{\mathrm{bulk}}$ denote the atomic energies in the surface and bulk environments, respectively. The LSE represents the energy loss caused by a single atom in a single-element (or unary) bulk environment when exposed to a single- or multi-element surface. This definition enables the quantification of the surface stability, even for surfaces in complex environments and multicomponent systems. The atomic energies in Equation (2) can be evaluated via energy density analysis (EDA) from first-principles calculations, such as DFT54, 55, 56, 57, 58, which was introduced by Nakai in 2002 54. EDA is accurate because it is based on first-principles calculations; however, it is not suitable for exhaustive calculations, such as those in the present study, because of its high computational cost. To reduce the computational cost, all the LSE values in this study were evaluated using a machine learning interatomic potential (MLIP), specifically the universal NNP M3GNet. In our previous study, we demonstrated that the intermediate information from M3GNet can efficiently and accurately represent the local environments of atoms in molecules.59 In the Behler–Parrinello NNP framework, the total energy $E_{\mathrm{tot}}$ of a system comprising $N$ atoms is calculated as the sum of the energies of the atoms.
@@
Yoo et al. demonstrated the ability to map the atomic energies obtained by NNPs onto NP and surface systems 60. Deringer et al. utilized the Gaussian approximation potential model to compute the atomic energies, which were then used to explore the configurational space and investigate the nature of defects in crystals 61, 62. MLIPs such as NNPs enable efficient evaluation of LSEs because the atomic energies of all adsorption sites in one system can be obtained in a single calculation.
@@ 2.3 Prediction model based on monometallic data
We introduce a predictive model for the adsorption energy of a molecule on a multi-element surface. This model is defined as a regression between the molecular adsorption energy on monometallic surfaces for each constituent element $\mathrm{M}$ of the multimetallic system and the LSE of the adsorption site prior to molecular adsorption. As the simplest model, we adopted the least-squares linear regression model expressed as follows:
@@
Here, $\alpha_{\mathrm{M}}$ and $\beta_{\mathrm{M}}$ denote the regression coefficients and constants, respectively, for each element $\mathrm{M}$. Simple regression makes the model explainable. $\alpha_{\mathrm{M}}$ represents the magnitude of the adsorption energy response to a change in LSE. $\beta_{\mathrm{M}}$ represents the adsorption energy of a molecule when the LSE is 0, that is, when the surface atom has the same energy as that in the bulk environment. For prediction, the adsorption energy is estimated by substituting the LSE values of the multi-element alloy surface prior to molecular adsorption into Equation (4).
@@ 3.1 Correlation Between Adsorption Energy and LSE
Fig.  2: Adsorption energies of CO calculated using the NNP for each on-top adsorption site of monometallic NPs $\mathrm{M}_{\mathrm{n}}$ with respect to the LSE. Solid lines represent the linear regressions of the adsorption energies of a CO molecule at the on-top sites of each monometallic NP according to Equation (4). Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  3: Adsorption energies of CO calculated using the NNP for each on-top adsorption site of HEA NPs with respect to the LSE. The inset presents a comparison of the distribution of adsorption energies for CO between the HEA NPs and monometallic NPs, with the x- and y-axes rescaled from Figure 2 for consistency. Dashed lines represent the linear regressions of the adsorption energies of a CO molecule at the on-top sites of each monometallic NP based on Equation (4). Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  4: Bar graphs in (a), (b), and (c) show the evaluated adsorption energies of CO on the (111) plane of a randomly selected $\text{HEA}_{201}$ at each on-top site of the facet, edge, and corner, respectively. The atoms of the selected plane are numbered as shown in (d). For each adsorption site, from left to right, the bar graph represents the adsorption energy obtained via direct evaluation using NNP (NNP direct), LSE-based prediction (LSE predict.), DFT, and LSE-based prediction parameterized by DFT data (I-LSE predict.). (d) Color mapping of the LSE values of the atoms on the selected plane and corresponding adsorption energies shown in (a), (b), and (c), along with the RMSE values relative to the DFT results. The asterisk indicates that the results of the structural optimization converge on the bridge site rather than the on-top site.
@@
Fig.  5: (a)Distributions of 122,000 LSE values of all the topmost layer atoms of the 1000 structure-optimized $\mathrm{HEA}_{201}$ NPs for each element and their sum, respectively. (b) Distributions of adsorption energies of CO for all the on-top adsorption sites predicted using the LSEs in (a) and Equation (4).
@@
The correlation between the LSE and the CO adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ of the monometallic NPs obtained using the NNP is shown in Figure 2. The solid lines represent linear regressions of the adsorption energies of a CO molecule at the top sites of each monometallic system. For all the metal elements, the relationship between the adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ of the CO molecule at the top site and the LSE is linear. The LSE values of all the adsorption sites for all elements range from approximately 0.2 to 1.2 eV. In other words, in all cases, the atomic energies are more unstable in the surface environment than in the bulk environment, which is reasonable given the lower coordination number in the surface environment. The adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ ranges from approximately -1.2 to -2.0 eV. Next, we examined the adsorption sites on the NPs. For all elements, the adsorption energy decreased in the following order: facets, edges, and corners. In contrast, the LSE values increased in the order of facets, edges, and corners. The LSE values and adsorption energies were concentrated at the facets, edges, and corners, and energy gaps existed between each group of adsorption sites. The $\mathrm{RMSE}$ was 0.035 eV, revealing a strong correlation between the LSE and the adsorption energy (see Figure 6 in the Appendix).
@@
Figure 3 presents the adsorption energies calculated directly using the NNP and their predicted values (dashed lines). The RMSE was 0.150 eV, which exceeded that of the unitary system. However, a strong correlation was observed between the LSE and adsorption energy in the HEA, indicating that the adsorption energy in a multicomponent environment can be effectively predicted (Figures 8(a) and (b) in the Appendix). This finding suggests that the LSE can efficiently and accurately predict the adsorption energies not only for unitary systems but also for complex systems such as HEAs. Notably, when we used the adsorption-energy range of -2.0 to -1.2 eV in the unitary system as the interpolation region for ${E_{\mathrm{ad}}(\mathrm{Predict.})}$, the predictions were more reliable than those for other ranges. In the extrapolation region of ${E_{\mathrm{ad}}(\mathrm{Predict.})}$, the difference from ${E_{\mathrm{ad}}(\mathrm{NNP})}$ increased, and a maximum shift of approximately 0.5 eV was observed. Next, we explored the trends for each elemental species. Figures 3 and  8(a) and (b) in the Appendix show that for all elemental species, the adsorption energies are nonlinearly estimated toward the unstable adsorption energy side. Additionally, the nonlinear region was dominated by adsorption at the facet sites. Compared to sites with low coordination numbers, such as corner and edge sites, the atoms on the facets were coordinated with eight or nine atoms. This increased coordination number renders them more sensitive to the surrounding environment than unitary systems. Consequently, the complex environment of HEAs may introduce unexpected nonlinearity into predictions. A potential solution to this problem is presented in the Discussion section.
@@
To show that this nonlinear trend can be captured by training the adsorption energies on HEA NPs, we applied kernel ridge regression (KRR) and quantum circuit learning (QCL) regression 63 to construct an adsorption energy prediction model for each elemental species at the HEA NP adsorption sites. Figures 8(c) and (d) in the Appendix show the correlation plots between the predicted and actual adsorption energies of CO on 14 HEA NP patterns generated by the regression model using KRR and QCL regression, respectively. The RMSEs of the adsorption energy predictions for all the adsorption sites provided by the KRR and QCL regression models were 0.0580 and 0.0579 eV, respectively, indicating comparable nonlinear regression performance. Compared with the uncorrected predictions based on the LSE, the RMSE values were reduced by a factor of approximately three.

### v3

@@
[Uncaptioned image] Computational screening is indispensable for the efficient design of high-entropy alloys (HEAs), which hold considerable potential for catalytic applications. However, the chemical space of HEAs is exponentially vast with respect to the number of constituent elements, making even machine learning-based screening calculations time-intensive. To address this challenge, we propose a rapid method for predicting HEA properties using data from monometallic systems (or few-component alloys). Central to our approach is the newly introduced local surface energy (LSE) descriptor, which captures local surface reactivity at atomic resolution. We established a correlation between LSE and adsorption energies using monometallic systems. Using this correlation in a linear regression model, we successfully estimated molecular adsorption energies on HEAs with significantly higher accuracy than a conventional descriptor (i.e., generalized coordination numbers). Furthermore, we developed high-precision models by employing both classical and quantum machine learning. Our method enabled CO adsorption-energy calculations for 1000 quinary nanoparticles, comprising 201 atoms each, within a few days, considerably faster than density functional theory, which would require hundreds of years or neural network potentials, which would have taken hundreds of days. The proposed approach accelerates the exploration of the vast HEA chemical space, facilitating the design of novel catalysts.
@@ 1 Introduction
Fig.  1: (a) Relationship between the molecular adsorption energy and the newly introduced descriptor, i.e., the local surface energy (LSE), for monometallic and high-entropy alloy (HEA) nanoparticle (NP) systems. The LSE descriptor developed in this study effectively captures the variation in adsorption energy across different adsorption sites. (b) Near-exponential increase in the number of NPs and adsorption sites as the number of elements in the NPs increases from monometallic to multicomponent systems.(c) The model is constructed through linear regression between the adsorption energies and the LSE using atomic energies from the NNP. The CO adsorption energies are evaluated on the ontop irreducible sites of facets (circles), edges (squares), and corners (triangles) of the monometallic NPs indicated by the blue symbols. (d) Fast catalytic property prediction workflow using neural network potentials (NNPs). The adsorption energy ${E_{\text{ad}}}$ is predicted from the LSE of the HEA NP surface prior to adsorption.
@@
To circumvent these limitations, neural network potentials (NNPs) based on the Behler–Parrinello framework 25, 26, 27, 28, 29 and graph neural networks 30, 31, 32, 33, offer promising solutions. Universal NNPs can encompass extensive elemental diversity and achieve high computational efficiency while maintaining accuracy on par with that of DFT 31, 32, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47. Recently, NNPs specializing in HEAs have emerged, made more lightweight through knowledge distillation 48. These advances have accelerated the prediction of catalytic properties; however, computational challenges remain.
In contrast, descriptor-based machine learning models offer scalability by predicting adsorption energies through generalized coordination numbers (GCNs), d-band centers, surface microstructural features, and local atomic environments, bypassing direct energy assessments 49, 50, 51, 52, 11, 53, 54, 55, 56, 10. These models have been proposed for predicting the adsorption energies of the remaining candidates by regressing the adsorption energies obtained from several first-principles calculations. Nonetheless, their applicability to HEAs is hampered by the complexity of alloy compositions and their dependence on extensive first-principles calculations.
@@ 2.1 Target systems and problems
We made the following assumptions regarding the structure and composition of HEA NPs: the structure is a truncated octahedron NP with 201 atoms in a face-centered cubic (fcc) arrangement (see 1(d)), the elemental composition ratio is as uniform as possible, and the atomic arrangement is randomly determined following a uniform distribution. We also assumed that the molecule occupied only a single top site. In HEA201, there are 122 sites. Even with these assumptions, the number of adsorption sites on HEA NPs is approximately $1.3\times 10^{138}$ (Figure 1(a)).
As a specific demonstration system, we measured the adsorption energies of CO molecules on IrPdPtRhRu NPs. In 2020, Wu et al. successfully synthesized IrPdPtRhRu NPs with nearly identical experimental composition ratios  6. The investigation of the adsorption characteristics of CO molecules is useful for evaluating the catalytic properties of HEAs 11, 57. As precursor systems for developing predictive models for HEAs, the monometallic NPs are of the truncated octahedron type corresponding to $\mathrm{M}_{\mathrm{n}}$ n=38, 79, 116, 201. For the on-top adsorption of the CO molecule on monometallic NPs $\mathrm{M}_{201}$, only the irreducible adsorption sites are calculated as shown in Figure 1(c). The adsorption energy $E_{\mathrm{ad}}$ is calculated as follows:
@@ 2.2 Local surface energy (LSE) descriptor
Fig.  2: Framework for predicting Local Surface Energies (LSEs) using a pre-trained universal neural network potential (NNP). The gray arrow indicates the computational flow from the pre-trained universal NNP, and the blue arrow represents the flow that generates LSEs. Atomic coordinates $\{Z_{I},R_{I}\}$ are processed by a graph neural network (GNN) layer to extract atomic features. A multi-Layer perceptron (MLP) layer predicts the atomic energies $E_{\text{atom}}^{I}$, which are reused to compute LSEs as the difference between surface atomic energies $E_{\text{atom}}^{\text{surf}}$ and bulk atomic energies $E_{\text{atom}}^{\text{bulk}}(Z_{\text{surf}})$. The resulting LSEs are visualized with a color map, highlighting local reactivity on the surface.
@@
where $E_{\mathrm{atom}}^{\mathrm{surf}}$ and $E_{\mathrm{atom}}^{\mathrm{bulk}}$ denote the atomic energies in the surface and bulk environments, respectively. The LSE represents the energy loss caused by a single atom in a single-element (or unary) bulk environment when exposed to a single- or multi-element surface. This definition enables the quantification of the surface stability, even for surfaces in complex environments and multicomponent systems. The atomic energies in Equation (2) can be evaluated via energy density analysis (EDA) from first-principles calculations, such as DFT58, 59, 60, 61, 62, which was introduced by Nakai in 2002 58. EDA is accurate because it is based on first-principles calculations; however, it is not suitable for exhaustive calculations, such as those in the present study, because of its high computational cost. To reduce the computational cost, all the LSE values in this study were evaluated using a machine learning interatomic potential (MLIP), specifically the universal NNP M3GNet. In our previous study, we demonstrated that the intermediate information from M3GNet can efficiently and accurately represent the local environments of atoms in molecules.63 In the Behler–Parrinello NNP framework, the total energy $E_{\mathrm{tot}}$ of a system comprising $N$ atoms is calculated as the sum of the energies of the atoms.
@@
Yoo et al. demonstrated the ability to map the atomic energies obtained by NNPs onto NP and surface systems 64. Deringer et al. utilized the Gaussian approximation potential model to compute the atomic energies, which were then used to explore the configurational space and investigate the nature of defects in crystals 65, 66. MLIPs such as NNPs enable efficient evaluation of LSEs because the atomic energies of all adsorption sites in one system can be obtained in a single calculation. Figure 2 illustrates the workflow for evaluating the Local Surface Energy (LSE) using a pre-trained universal NNP.
@@ 2.3 Prediction model based on monometallic data
We introduce a predictive model for the adsorption energy of a molecule on a multi-element surface. This model is defined as a regression between the molecular adsorption energy on monometallic surfaces for each constituent element $\mathrm{M}$ of the multimetallic system and the LSE of the adsorption site prior to molecular adsorption. Figure 1(c) illustrates the workflow for constructing the predictive model. As the simplest model, we adopted the least-squares linear regression model expressed as follows:
@@
Here, $\alpha_{\mathrm{M}}$ and $\beta_{\mathrm{M}}$ denote the regression coefficients and constants, respectively, for each element $\mathrm{M}$. Simple regression makes the model explainable. $\alpha_{\mathrm{M}}$ represents the magnitude of the adsorption energy response to a change in LSE. $\beta_{\mathrm{M}}$ represents the adsorption energy of a molecule when the LSE is 0, that is, when the surface atom has the same energy as that in the bulk environment. For prediction, the adsorption energy is estimated by substituting the LSE values of the multi-element alloy surface prior to molecular adsorption into Equation (4). Figure 1(d) illustrates the workflow for predicting CO adsorption energies on HEA NPs using the proposed model. Here, the prediction model using adsorption energies from the NNP for regression is referred to as the LSE-based prediction model, while the model using adsorption energies from DFT is referred to as the Improved-LSE (I-LSE) prediction model. Details of the model construction can be found in the Computational details section.
Fig.  3: Adsorption energies of CO calculated using the NNP for each on-top adsorption site of monometallic NPs $\mathrm{M}_{\mathrm{n}}$ with respect to the LSE. Solid lines represent the linear regressions of the adsorption energies of a CO molecule at the on-top sites of each monometallic NP according to Equation (4). Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  4: Adsorption energies of CO calculated using the NNP for each on-top adsorption site of 20 HEA201 NPs with respect to the LSE. The inset presents a comparison of the distribution of adsorption energies for CO between the HEA NPs and monometallic NPs, with the x- and y-axes rescaled from Figure 3 for consistency. Dashed lines represent the linear regressions of the adsorption energies of a CO molecule at the on-top sites of each monometallic NP based on Equation (4). Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  5: Bar graphs in (a), (b), and (c) show the evaluated adsorption energies of CO on the (111) plane of a randomly selected $\text{HEA}_{201}$ at each on-top site of the facet, edge, and corner, respectively. The atoms of the selected plane are numbered as shown in (d). For each adsorption site, from left to right, the bar graph represents the adsorption energy obtained via direct evaluation using NNP (NNP direct), LSE-based prediction (LSE predict.), DFT, and LSE-based prediction parameterized by DFT data (I-LSE predict.). (d) Color mapping of the LSE values of the atoms on the selected plane and corresponding adsorption energies shown in (a), (b), and (c), along with the RMSE values relative to the DFT results. The asterisk indicates that the results of the structural optimization converge on the bridge site rather than the on-top site.
@@ 3.1 Correlation Between Adsorption Energy and LSE
Fig.  6: (a)Distributions of 122,000 LSE values of all the topmost layer atoms of the 1000 structure-optimized $\mathrm{HEA}_{201}$ NPs for each element and their sum, respectively. (b) Distributions of adsorption energies of CO for all the on-top adsorption sites predicted using the LSEs in (a) and Equation (4). Solid lines represent the values for monometallic NP M201 (M=Ir, Pd, Pt, Rh, and Ru).
@@
The correlation between the LSE and the CO adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ of the monometallic NPs obtained using the NNP is shown in Figure 3. The solid lines represent linear regressions of the adsorption energies of a CO molecule at the top sites of each monometallic system. For all the metal elements, the relationship between the adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ of the CO molecule at the top site and the LSE is linear. The LSE values of all the adsorption sites for all elements range from approximately 0.2 to 1.2 eV. In other words, in all cases, the atomic energies are more unstable in the surface environment than in the bulk environment, which is reasonable given the lower coordination number in the surface environment. The adsorption energy $E_{\mathrm{ad}}(\mathrm{NNP})$ ranges from approximately -1.2 to -2.0 eV. Next, we examined the adsorption sites on the NPs. For all elements, the adsorption energy decreased in the following order: facets, edges, and corners. In contrast, the LSE values increased in the order of facets, edges, and corners. The LSE values and adsorption energies were concentrated at the facets, edges, and corners, and energy gaps existed between each group of adsorption sites. The $\mathrm{RMSE}$ was 0.035 eV, revealing a strong correlation between the LSE and the adsorption energy (see Figure 9 in the Details of the LSE-based regression models in Appendix).
@@
Figure 4 presents the adsorption energies calculated directly using the NNP and their predicted values (dashed lines). The RMSE was 0.150 eV, which exceeded that of the unitary system. However, a strong correlation was observed between the LSE and adsorption energy in the HEA, indicating that the adsorption energy in a multicomponent environment can be effectively predicted (Figures 11(a) and (b) in the Appendix). This finding suggests that the LSE can efficiently and accurately predict the adsorption energies not only for unitary systems but also for complex systems such as HEAs. Notably, when we used the adsorption-energy range of -2.0 to -1.2 eV in the unitary system as the interpolation region for ${E_{\mathrm{ad}}(\mathrm{Predict.})}$, the predictions were more reliable than those for other ranges. In the extrapolation region of ${E_{\mathrm{ad}}(\mathrm{Predict.})}$, the difference from ${E_{\mathrm{ad}}(\mathrm{NNP})}$ increased, and a maximum shift of approximately 0.5 eV was observed.
@@
Next, we explored the trends for each elemental species. Figures 4 and  11(a) and (b) in the Appendix show that for all elemental species, the adsorption energies are nonlinearly estimated toward the unstable adsorption energy side. Additionally, the nonlinear region was dominated by adsorption at the facet sites. Pd atoms in HEA nanoparticles clearly show this non-linear behavior; unlike Pt atoms, they show a reversed pattern of adsorption energies. Compared to sites with low coordination numbers, such as corner and edge sites, the atoms on the facets were coordinated with eight or nine atoms. This increased coordination number renders them more sensitive to the surrounding environment than unitary systems. Consequently, the complex environment of HEAs may introduce unexpected nonlinearity into predictions. To show that this nonlinear trend can be captured by training the adsorption energies on HEA NPs, we applied kernel ridge regression (KRR) and quantum circuit learning (QCL) regression 67 to construct an adsorption energy prediction model for each elemental species at the HEA NP adsorption sites. Figures 11(c) and (d) in the Appendix show the correlation plots between the predicted and actual adsorption energies of CO on 14 HEA NP patterns generated by the regression model using KRR and QCL regression, respectively. The RMSEs of the adsorption energy predictions for all the adsorption sites provided by the KRR and QCL regression models were 0.0580 and 0.0579 eV, respectively, indicating comparable nonlinear regression performance. Compared with the uncorrected predictions based on the LSE, the RMSE values were reduced by a factor of approximately three.